If x= a cos^3A, y=a sin^3 A then 
a) x^2/3 +y^2/3= a^2/3
B) x^2+y^2=a^2
c) y^2 =4ax

Grade:11

1 Answers

Vikas TU
7 years ago
Dear Nidhi,
x= a cos^3A
then cos A = (x/a)^(1/3)......................(1)
y=a sin^3 A
sin A = (y/a)^(1/3)..................................(2)
 
squaring eqns. (1)  and (2)
(sin^2A + cos^2A) =  (x/a)^(2/3) + (y/a)^(2/3)
 (x/a)^(2/3) + (y/a)^(2/3) = 1
x^2/3 + y^2/3 = a^2/3
 
option A
plz approve!